Circular No. 6969

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bw IN ESO 184-G82 L. E. Kay, J. P. Halpern, and K. M. Leighly, Columbia University;

S. Heathcote, Cerro Tololo Interamerican Observatory (CTIO); and A. M. Magalhaes, Astronômico e Geofísico, Universidade de São Paulo, report on spectropolarimetry of SN 1998bw, possibly associated with GRB 980425 (cf. IAUC 6901, 6903), obtained with the CTIO 4-m telescope on June 20, 21, and 23 UT: “Over the wavelength range 390–750 nm, we measure intrinsic linear polarization of 0.53 ± 0.08 percent at position angle 49^{o} ± 3^{o}, after correcting for Galactic interstellar polarization using the star HD 184100, which has polarization of 0.75 ± 0.01 percent at p.a. 176^{o}.5 ± 2^{o}.5. This measured interstellar polarization is consistent with the Galactic extinction in this direction, estimated to be E(B − V) = 0.059 from IRAS maps, or E(B − V) = 0.079 from 21-cm H I. Interstellar polarization in the host galaxy ESO 184-G82 is expected to be negligible based on the relative absence of Na I D absorption at z = 0.00841 ± 0.00005, the redshift of the environment of the supernova from narrrow H II region emission lines in its spectrum. Polarization appears highest in between emission features in the total flux spectrum, which strengthens the interpretation of the polarization as intrinsic to the supernova. This modest polarization is less than that of some type-II supernovae, but greater than that of type-Ia supernovae, which are generally unpolarized. This supports the interpretation of SN 1998bw, a peculiar type-Ic supernova, as a core-collapse event in which the observed polarization is due to moderate asymmetry in either the photosphere of the ejecta or an overlying scattering envelope. However, this result does not strongly constrain arguments about whether some supernovae emit γ-ray bursts, since such emission may come from a mildly relativistic shock associated with the radio emission and above the optical photosphere, without any requirements on beaming or orientation.”

A. V. Filippenko, University of California at Berkeley, comments on the total flux spectrum obtained above: “The spectrum most closely resembles those of the peculiar SN 1997ef, but perhaps evolving more slowly. It is not typical of type-Ic supernovae; indeed, the spectrum does not match any of the known spectral classes, but perhaps ‘peculiar type Ic’ is the best choice at this time. The object appears to be starting a transition to the nebular phase, but identifications of the broad emission features are uncertain. The feature at 650 nm is probably [Fe III], as is the one at 455 nm (perhaps with some contribution from Mg I]). That at 590 nm may be a mixture of [Co III] and Na D. The strongest feature, at 545 nm, may consist of Fe, Co, and Ti lines.”

1998 July 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green